OpenStudy (anonymous):

Find the exact value of the expression given that secx=3/2, cscy=3 and x and y are in quadrant 1

OpenStudy (anonymous):

Oh and the expression is sin (x+y)

OpenStudy (anonymous):

you need a bunch of numbers you need \(\sin(x), \cos(x), \sin(y), \cos(y)\) do yuou know how to find them ?

OpenStudy (anonymous):

I know how to find the cos x and siny but not the others

OpenStudy (anonymous):

\[\cos(x)=\frac{2}{3}\]and you need \(\sin(x)\) draw a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

the third side, by pythagoras , is \[\sqrt{3^2-2^2}=\sqrt5\]

OpenStudy (anonymous):

OOPS! \[\huge \sin(x)=\frac{\sqrt5}{3}\] sorry

OpenStudy (anonymous):

Ok that makes sense. And you do that similarly to cosy, right? Is cosy=(2sqrt2)/3

OpenStudy (anonymous):

yes you do idk i didn't do it but i can check real quick

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

Ok well if that is right I should be able to do the rest. Thank you so much!

OpenStudy (anonymous):

yw, "addition angle" formula from here on in plug in the numbers, see what you get